OpenStudy (anonymous):

How do you do 1.5/2.5 of 15?

OpenStudy (phi):

"of" means multiply \[ \frac{1.5}{2.5} \cdot 15 \] to get rid of the decimals, you can multiply top and bottom by 10 \[ \frac{15}{25} \cdot 15 \] can you finish?

OpenStudy (phi):

I would first simplify the fraction by dividing top and bottom by 5 what do you get ?

OpenStudy (anonymous):

like what is 1.5/2.5 of 15

OpenStudy (phi):

re-read my post more carefully.

OpenStudy (phi):

do you know how to simplify 15/25 by dividing the top number by 5 and the bottom number by 5?

OpenStudy (anonymous):

you get 3/5

OpenStudy (phi):

yes, that means the problem is \[ \frac{3}{5} \cdot 15 \] notice you can simplify more by dividing the bottom by 5 and the 15 (which is up top) by 5

OpenStudy (anonymous):

45/75

OpenStudy (anonymous):

The end answer is 9 but how do I get it?

OpenStudy (phi):

when you multiply fractions you multiply top times top and bottom times bottom if you have a whole number like 15, write it as 15/1 so it has a "bottom" but it is better to first simply first. you have \[ \frac{ 3}{5} \cdot \frac{15}{1} \\ \frac{3 \cdot 15}{5\cdot 1}\] I would not multiply 5*15 up top right away instead, divide the bottom by 5 and the top by 5 to get smaller numbers

OpenStudy (anonymous):

But im not multiplying

OpenStudy (anonymous):

Im trying to find out what is 1.5/2.5 out of 15

OpenStudy (anonymous):

and how to work it out

OpenStudy (anonymous):

the end answer is 9

OpenStudy (anonymous):

but how do I get it?

OpenStudy (phi):

English is confusing when describing math (which is why equations are easier to understand .. once you learn how) but you can get the idea this way: what is 1/2 of 2 ? notice the answer is 1. and 1/2 * 2 is also 1 It turns out the "of" means multiply

OpenStudy (anonymous):

Oh ok

OpenStudy (phi):

so try to simplify \[ \frac{3 \cdot 15}{5\cdot 1} \] by dividing top and bottom by 5

OpenStudy (anonymous):

3/3

OpenStudy (anonymous):

Cross multiply?

OpenStudy (phi):

in the bottom you have a 5. divide by 5 you get 1 up top you have a 15 what is 15/5 ?

OpenStudy (phi):

Once you learn the idea you will be able to do these problems

OpenStudy (anonymous):

3

OpenStudy (phi):

what we did was simplify \[ \frac{3 \cdot 15}{5\cdot 1} = \frac{3 \cdot 3}{1\cdot 1} \] now you can do the multiply and divide and get a simple number.

OpenStudy (anonymous):

im so confused

OpenStudy (anonymous):

How you get a one

OpenStudy (phi):

you could have done \[ \frac{3 \cdot 15}{5\cdot 1} = \frac{45}{5} \] and simplify to get 9 but it's easier for me to divide the 5 into the 15 right away

OpenStudy (phi):

if you have \[ \frac{15}{5} \] you can divide the top by 5 and the bottom by 5 to get \[ \frac{3}{1} \]

OpenStudy (anonymous):

Thats easier for me. Thank you

OpenStudy (phi):

maybe. but notice if you had \[ \frac{256}{32} \] and no calculator. It should be clear you can divide top and bottom by 2 so you could simplify the problem to \[ \frac{128}{16} \] and then divide top and bottom by 2 again \[ \frac{64}{8} \] and now you might notice 64/8 is 8

OpenStudy (phi):

but if you did not notice 64/8 is 8 you could simplify by dividing top and bottom by 2 to get 32/4 and then 16/2 and then 8/1 or just 8